### v1

@@ Abstract
Index Terms:
fermion-qubit mapping, quantum computing, ancilla qubits, quadratic assignment.
@@ I Introduction
Fermion-qubit mappings in the literature generally fall into two paradigms: ancilla-free mappings which use the minimum number of qubits to describe the fermionic algebra [9, 10, 11, 12, 13, 14, 15, 16] and local encodings which employ a number of ancilla qubits that scales linearly with the system size [9, 17, 18, 19, 20, 21, 22, 23, 24]. As seen in Figure 1, local encodings are designed to produce qubit Hamiltonians that are sums of local Pauli operators for fermionic systems such as the Fermi–Huubard model [25]. Meanwhile, ancilla-free mappings have been optimized to produce Pauli representations of fermionic hopping terms with logarithmically scaling support size through the Bravyi–Kitaev [9] and Ternary Tree transformations [12, 11]. However, in the near- and mid-term future for quantum computing, it is not necessarily the case that local encodings are always preferable to ancilla-free mappings [26, 27], with qubit number itself being a scarce resource [28] competing with the Pauli weights of the qubit Hamiltonian which serve as theoretical complexity parameters.
@@ III Background & Notation
In this section, we give a brief overview of fermionic systems and fermion-qubit mappings, defining the ‘linear encoding’ class of ancilla-free mappings and establishing the conditions for existence of mappings with ancilla qubits. We introduce a variety of notational conventions that will be used throughout the paper, and adopt zero-indexing, i.e. $[j]=\{0,1,\dots,j-1\}$.
@@ III-A Fermionic Systems vs Qubit Systems
A mathematical description of fermionic statistics requires choosing an ordering $i\in[n]$ for the modes: while physically unimportant, the order enables the description of the $(2^{n})$-dimensional Hilbert space of the system as $\mathcal{H}_{f}^{n}=\text{span}\{|{\mathbf{f}})\mid\mathbf{f}\in\mathds{Z}_{2}^{n}\}$, where the Fock basis state $|{\mathbf{f}})$ is the antisymmetrized wavefunction of an $n$–particle system with mode $i$ occupied if and only if $f_{i}=1$. The algebra of fermionic operators $B(\mathcal{H}_{f}^{n})$ derives from the creation and annihilation operators $a_{i}^{(\dagger)}$ for $i\in[n]$, which act on the Fock basis states via
$a_{i}^{(\dagger)}|{\mathbf{f}})=\begin{cases}(-1)^{\sum_{j<i}f_{j}}|{(\mathbf{f}+\mathbf{e}_{i})\bmod 2})&f_{i}=1\,(0)\\ 0&f_{i}=0\,(1)\,,\end{cases}$ (1)
@@ III-B Fermion-qubit mappings and the Jordan–Wigner transformation
$\mathfrak{m}_{\text{JW}}:\mathcal{H}_{f}^{n}\ni|{\mathbf{f}})\longmapsto\ket{f_{0}}\ket{f_{1}}\dots\ket{f_{n-1}}\in\mathcal{H}_{2}^{\otimes n}\,.$ (6)
@@ III-C Optimization problems for fermion-qubit mappings
Each hopping term maps to a sum of four Pauli operators, each with weight $|i{-}j|{+}1$, as in Figure 2. The number and interaction terms map to local sums of Pauli operators with weights 1 and 2, respectively, regardless of the total number of modes $n$. The presence of hopping terms means that overall, the Jordan–Wigner transformation produces terms of weights up to $n$ in simulation tasks for $H_{f}$. In contrast, other ancilla-free mappings such as the Bravyi–Kitaev [9] and Ternary Tree transformations [12, 11] produce qubit representations of the hopping terms with weights up to ${\sim}\log_{2}n$ and ${\sim}\log_{3}(2n)$. While this is an advantage over the Jordan–Wigner transformation in terms of total and maximum Pauli weight, we should note that the number and interaction terms do not transform into local qubit operators under these other mappings.
@@ III-D Majorana operators and linear encodings
$\mathfrak{m}_{U}:\mathcal{H}_{f}^{n}\ni|{\mathbf{f}})\longmapsto\ket{U\mathbf{f}}\in\mathcal{H}_{2}^{\otimes n}$ (17)
@@
$\mathfrak{m}_{\text{BK}}:|{\mathbf{f}})$ |$\longmapsto\ket{U_{\text{BK}}\mathbf{f}}\,,$ (20)
@@
$\mathfrak{m}_{\text{TT}}:|{\mathbf{f}})\longmapsto\ket{U_{\text{TT}}\mathbf{f}}\,,$ (22)
@@ III-E Using ancilla qubits in fermion-qubit mappings
    where the $(2^{n})$-dimensional Hilbert space $\mathcal{H}$ is the space stabilized by any qubit operator in the kernel of $\mathfrak{m}^{\dagger}$. That is, if the qubit operator $A\in B(\mathcal{H}_{2}^{\otimes m})$ is an image of the fermionic identity in the sense that $\mathfrak{m}^{\dagger}(A)=1$, then $A\ket{\psi}=\ket{\psi}$ for all |$\ket{\psi}\in\mathcal{H}$.
@@ IV Our Methodology
The second approach, in Section IV-B, introduces a new strategy to incrementally add ancilla qubits to the Jordan–Wigner transformation in order to produce a qubit Hamiltonian with lower maximum and total Pauli weights than the mappings with optimal ordering from Section IV-A. This method generalises the approach in [29] to add more than two qubits, and applies to arbitrary fermionic Hamiltonian graphs rather than just the square lattice.
@@
[…]
@@ IV-A Fermionic Order as Quadratic Assignment
where $L_{ij}=1$ if $i<j$ and 00 otherwise. Each column of $U$ and each row of $F,P,R$ then correspond to the set families of the same name in (18), e.g. $U_{\ast j}=U(j)$ and $F_{i}=F(i)$.
Using (34) and (35), we may use the matrix definitions of $F$, $P$ and $R$ to define a matrix $\mathcal{D}\in\mathbb{N}^{n\times n}$ representing the total (=$\oplus$=+⊕ = +) or maximum ($\oplus=\max$) Pauli weight of $H_{q}(\sigma)$ consisting of the elementwise $\oplus$-sum of the weight of number terms, hopping terms with a nontrivial real part, hopping terms with a nontrivial complex part, and interaction terms:
@@ IV-B Incremental addition of ancilla qubits
where $\sigma:V\rightarrow[n]$ is the optimal fermionic label order for total or maximum Pauli weight from Section IV-A, depending on the cost function of choice. Unlike in (12), Equation (39) conceals the linear combination of the two-qubit operators $X_{i}X_{j}$, $Y_{i}Y_{j}$, $X_{i}Y_{j}$ and $Y_{i}X_{j}$ as $(...)_{ij}$ for brevity. The unwritten terms in (39) are the the number and interaction terms, which are local under the Jordan–Wigner transformation. As per statement 2, the terms of $H_{q}(\sigma)$ have the same Hermiticity and commutation relations relative to the terms in $H_{f}$.
Figure 7: Introducing an ancilla qubit allows modification of the hopping terms in a Jordan–Wigner Hamiltonian to reduce the Pauli weight of some terms.
We wish to construct an $(n{+}1)$-qubit Hamiltonian $H_{q}^{\prime}$ that also satisfies statement 2, while reducing the Pauli weights of some of the hopping terms. Define $H_{q}^{\prime}$ to be an $(n{+}1)$-qubit Hamiltonian with identical number and interaction terms to $H_{q}(\sigma)$, but with each hopping term modified via
@@
The question is then, how much does $\kappa_{ij}$ affect the Pauli weight of a Jordan–Wigner hopping term? For Option 2, this is fairly trivial: the weight of an $(i,j)$ hopping term in the new Hamiltonian $H_{q}^{\prime}$ increases by |$|\{i,j\}\cap\mathcal{I}|\bmod 2$, corresponding to the case where a nontrivial gate is placed on the ancilla.
@@
Subsequent ancilla qubits can be added by renaming $H_{q}^{\prime}$ to $H_{q}(\sigma)$ and iteratively applying the procedure described above. In introducing the $k$th ancilla qubit, choose a qubit subset ${\mathcal{I}_{k}}$ and define $\kappa_{ij}$ as in (41), with $Z_{\mathcal{I}_{k}}$ acting on the first $n$ qubits and the single-qubit Pauli acting on the $k$th ancilla qubit, leaving the previous ancilla qubits untouched. Perform the conditional modification of the hopping terms via Option 1 and Option 2 to produce an $(n{+}k)$-qubit Hamiltonian with lower Pauli weights. Continue to add ancilla qubits incrementally until exhausting the available supply, or if there is no further reduction in the total or maximum Pauli weight of the hopping terms in the modified qubit Hamiltonian.
@@ IV-C Constant-ancilla method for other mappings
Our method for constructing constant-ancilla fermion-qubit mappings in Section IV-B begins by using the Jordan–Wigner transformation to generate the base ancilla-free Hamiltonian in (39). If we had instead used a different linear encoding, we predict that it would generally be computationally hard to devise an effective set of commuting Pauli strings $P_{0},P_{1},\dots,P_{p-1}$ for incremental ancilla addition because number, hopping and interaction terms map to Pauli strings with less obvious nonlocal structure than the $Z_{i+1}\dots Z_{j-1}$ factors of the Jordan–Wigner transformations. In modifying the Hamiltonian terms from $H_{q}(\sigma)$ to $H_{q}^{\prime}$ with the addition of the $k$th ancilla qubit, the protocol would need to search for Pauli operators $P_{k}=X_{\mathcal{I}_{k}}Y_{\mathcal{J}_{k}}Z_{\mathcal{K}_{k}}$ such that multiplying the terms of $H_{q}(\sigma)$ by $P_{k}$ reduces the total or maximum Pauli weight. As discussed elsewhere [19, 29] a new requirement emerges that $[P_{j},P_{k}]=0$ for all $j\in[k]$ in order to satisfy statement 2. Again, the reason this was straightforward for the Jordan–Wigner case was because we knew a priori that the problematic parts of the hopping terms would consist only of $Z$ matrices. Nevertheless, we see this as a worthwhile line of inquiry because of the potential for other mappings to reduce the weight of qubit Hamiltonians from the outset.
TABLE I: A comparison of the Pauli weight for Jordan–Wigner, Bravyi–Kitaev, Parity Basis and Ternary Tree mappings of fermionic systems with Hamiltonian graphs that are square grids of size 3$\times$3 to 15$\times$15. Orderings are optimized for total Pauli weight, max Pauli weight, and total Pauli weight assuming real valued coefficients and the absence of operator terms, corresponding to the Fermi–Hubbard model.
@@
[…]
@@
TT 313 628 1080 1676 2375 3237 4303 5473 6799 8342 9853 11844 13942
[…]
@@
64-mode system graph JW TT
@@ V Experiments
Figure 8: Comparison of the total Pauli weight of hopping terms ($\mathcal{D}=\text{ReHop}+\text{ImHop}$) corresponding to Jordan–Wigner with an increasing number of ancilla qubits against Bravyi–Kitaev (8) and Ternary Tree (8) for the 8$\times$8 and 5$\times$8 grid respectively.
V-A Encoding comparison with optimal fermionic order
In this section, we examine the difference in the total and maximum Pauli weights yielded by the four primary encodings. In Table I, we compare the results of the Jordan–Wigner, Bravyi–Kitaev, Parity Basis and Ternary Tree transformations applied to a series of square grids of increasing size. We show results for three different optimization objectives: total Pauli weight, maximum Pauli weight, and total Pauli weight for a model with only real-valued coefficients and no operator terms, which corresponds to the popular Fermi–Hubbard model [41]. Here, we see that Jordan–Wigner is the preferred encoding for smaller grids, allowing orderings which produce a small maximum and total Pauli weight compared to other methods. However, for all but the smallest input, Ternary Tree remains comparable, and eventually emerges as the superior ordering. Bravy–Kitaev’s better scaling similarly allows it to outperform Jordan–Wigner for larger grids.
@@ V-B Jordan–Wigner transformations with optimal ancilla use
In Figure 8, we compare Jordan–Wigner with multiple ancilla qubits with the ancilla-free Bravyi–Kitaev and Ternary Tree transformations on the 8$\times$8 and 5$\times$8 grids, respectively. These graphs are chosen so that the traditional definitions from (21) and (23) hold, respectively. Here, we see that adding only a few ancilla qubits can significantly improve the Pauli weights associated with the Jordan–Wigner transformation, to the point where it overtakes the Pauli weight of Ternary Tree. In the case of the 8$\times$8 grid, using 10 ancilla qubits allows us to reduce the Pauli weight of the hopping terms in the final Jordan–Wigner Hamiltonian by 39.3% compared to the ancilla-free Jordan–Wigner Hamiltonian.
@@
64-mode system graph JW TT JW + 10 ancillas
@@ VI Conclusion
We have introduced a cost model for the total and maximum Pauli weight of the output of linear encodings, a type of ancilla-free fermion-qubit mapping, and have found optimal fermionic orderings for several popular transformations using quadratic assignment. Moreover, we find an optimal strategy to cancel out disjoint strings of Pauli-$Z$ gates, utilizing ancilla qubits to reduce the Pauli weight of qubit Hamiltonians associated with the Jordan–Wigner transformation. According to our results, the latter strategy requires only of 10 ancilla qubits to reduce Hamiltonian Pauli weight by as much as 67%, outperforming the optimal ancilla-free Ternary Tree transformation.

### v2

@@ I Introduction
Fermion-qubit mappings in the literature generally fall into two paradigms: ancilla-free mappings which use the minimum number of qubits to describe the fermionic algebra [4, 34, 19, 40, 27, 26, 14, 10] and local encodings which employ a number of ancilla qubits that scales linearly with the system size [4, 39, 8, 38, 35, 12, 9, 1, 16]. As seen in Figure 1, local encodings are designed to produce qubit Hamiltonians that are sums of local Pauli operators for fermionic systems such as the Fermi–Hubbard model [17]. Meanwhile, ancilla-free mappings have been optimized to produce Pauli representations of fermionic hopping terms with logarithmically scaling support size through the Bravyi–Kitaev [4] and Ternary Tree transformations [40, 19]. However, in the near- and mid-term future for quantum computing, it is not necessarily the case that local encodings are always preferable to ancilla-free mappings [6, 29], with qubit number itself being a scarce resource [2] competing with the Pauli weights of the qubit Hamiltonian which serve as theoretical complexity parameters.
@@ III Background & Notation
In this section, we give a brief overview of fermionic systems and fermion-qubit mappings, defining the ‘linear encoding’ class of ancilla-free mappings and establishing the conditions for existence of mappings with ancilla qubits. We introduce a variety of notational conventions that will be used throughout the paper, and adopt zero-indexing, i.e. $j]=\{0,1,\dots,j-1\}$.
@@ III-A Fermionic Systems vs Qubit Systems
A mathematical description of fermionic statistics requires choosing an ordering $i\in[n]$ for the modes: while physically unimportant, the order enables the description of the $(2^{n})$-dimensional Hilbert space of the system as $\mathcal{H}_{f}^{n}=\text{span}\{|{\mathbf{f}})\mid\mathbf{f}\in\mathds{Z}_{2}^{n}\}$, where the Fock basis state $|{\mathbf{f}})$ is the antisymmetrized wavefunction of an nn–particle system with mode $i$ occupied if and only if $f_{i}=1$. The algebra of fermionic operators $B(\mathcal{H}_{f}^{n})$ derives from the creation and annihilation operators $a_{i}^{(\dagger)}$ for $i\in[n]$, which act on the Fock basis states via
[…]
@@ III-B Fermion-qubit mappings and the Jordan–Wigner transformation
[…]
@@ III-C Optimization problems for fermion-qubit mappings
Each hopping term maps to a sum of four Pauli operators, each with weight |ij|+1|i{-}j|{+}1, as in Figure 2. The number and interaction terms map to local sums of Pauli operators with weights 1 and 2, respectively, regardless of the total number of modes $n$. The presence of hopping terms means that overall, the Jordan–Wigner transformation produces terms of weights up to $n$ in simulation tasks for $H_{f}$. In contrast, other ancilla-free mappings such as the Bravyi–Kitaev [4] and Ternary Tree transformations [40, 19] produce qubit representations of the hopping terms with weights up to ${\sim}\log_{2}n$ and ${\sim}\log_{3}(2n)$. While this is an advantage over the Jordan–Wigner transformation in terms of total and maximum Pauli weight, we should note that the number and interaction terms do not transform into local qubit operators under these other mappings.
@@ III-D Majorana operators and linear encodings
[…]
@@
Here, the expression $U(i)$ denotes the row indices of nonzero elements in the $i$th column of $U$. If the expression $F(i)$ denotes the column indices of nonzero elements in the $i$th row of $U^{-1}$, then $P(i)=F(0)\,\triangle\,F(1)\,\triangle\,\dots\,\triangle\,F(i-1)$ and $R(i)=P(i)\,\triangle\,F(i)$, where $\triangle\,$ denotes the symmetric difference of two sets: $A\,\triangle\,B=(A\cup B)\backslash(A\cap B)$.
@@
[…]
@@
[…]
@@ III-E Using ancilla qubits in fermion-qubit mappings
[…]
@@ IV Our Methodology
The second approach, in Section IV-B, introduces a new strategy to incrementally add ancilla qubits to the Jordan–Wigner transformation in order to produce a qubit Hamiltonian with lower maximum and total Pauli weights than the mappings with optimized orderings from Section IV-A. This method generalises the approach in [11] to add more than two qubits, and applies to arbitrary fermionic Hamiltonian graphs rather than just the square lattice.
@@
$a_{i}^{(\dagger)}=\frac{1}{2}(\gamma_{i}\pm\mathrm{i}\mkern 1.0mu\overline{\gamma}_{i})\,,\quad a_{i}^{\dagger}a_{i}=\frac{1}{2}(1+\mathrm{i}\mkern 1.0mu\gamma_{i}\overline{\gamma}_{i})$ (30)
@@ IV-A Fermionic Order as Quadratic Assignment
where $L_{ij}=1$ if $i<j$ and 0 otherwise. Each column of $U$ and each row of $F,P,R$ then correspond to the set families of the same name in (18), e.g. $U_{\ast j}=U(j)$ and $F_{i}=F(i)$.
Using (34) and (35), we may use the matrix definitions of $F$, $P$ and $R$ to define a matrix $\mathcal{D}\in\mathbb{N}^{n\times n}$ representing the total ($\oplus=+$) or maximum ($\oplus=\max$) Pauli weight of $H_{q}(\sigma)$ consisting of the elementwise $\oplus$-sum of the weight of number terms, hopping terms with a nontrivial real part, hopping terms with a nontrivial complex part, and interaction terms:
@@ IV-B Incremental addition of ancilla qubits
where $\sigma:V\rightarrow[n]$ is an optimized fermionic label order for total or maximum Pauli weight from Section IV-A, depending on the cost function of choice. Unlike in (12), Equation (39) conceals the linear combination of the two-qubit operators $X_{i}X_{j}$, $Y_{i}Y_{j}$, $X_{i}Y_{j}$ and $Y_{i}X_{j}$ as $(...)_{ij}$ for brevity. The unwritten terms in (39) are the the number and interaction terms, which are local under the Jordan–Wigner transformation. As per statement 2, the terms of $H_{q}(\sigma)$ have the same Hermiticity and commutation relations relative to the terms in $H_{f}$.
Figure 7: One iteration of the method for incrementally adding ancilla qubits to Jordan–Wigner transformations and the corresponding reduction in Pauli weights of a qubit Hamiltonian. The ancilla comes with the user’s choice of Pauli string $Z_{\mathcal{I}}$: in constructing the Hamiltonian $H_{q}^{\prime}$, which is equivalent to $H_{q}(\sigma^{\prime})$, the user has the choice to modify each term of the original Hamiltonian by $Z_{\mathcal{I}}$ or leave it unchanged up to local action on the ancilla, leading to a reduction in Pauli weights.
We wish to construct an (n+1)(n{+}1)-qubit Hamiltonian $H_{q}^{\prime}$ that also satisfies statement 2, while reducing the Pauli weights of some of the hopping terms. Define $H_{q}^{\prime}$ to be an (n+1)(n{+}1)-qubit Hamiltonian with identical number and interaction terms to $H_{q}(\sigma)$, but with each hopping term modified via
@@
The question is then, how much does $\kappa_{ij}$ affect the Pauli weight of a Jordan–Wigner hopping term? For Option 2, this is fairly trivial: the weight of an $(i,j)$ hopping term in the new Hamiltonian $H_{q}^{\prime}$ increases by $|\{i,j\}\cap\mathcal{I}|\bmod 2$, corresponding to the case where a nontrivial gate is placed on the ancilla.
@@
Subsequent ancilla qubits can be added by renaming $H_{q}^{\prime}$ to $H_{q}(\sigma)$ and iteratively applying the procedure described above. In introducing the $k$th ancilla qubit, choose a qubit subset $\mathcal{I}_{k}$} and define $\kappa_{ij}$ as in (41), with $Z_{\mathcal{I}_{k}}$ acting on the first $n$ qubits and the single-qubit Pauli acting on the $k$th ancilla qubit, leaving the previous ancilla qubits untouched. Perform the conditional modification of the hopping terms via Option 1 and Option 2 to produce an (n+k)(n{+}k)-qubit Hamiltonian with lower Pauli weights. Continue to add ancilla qubits incrementally until exhausting the available supply, or if there is no further reduction in the total or maximum Pauli weight of the hopping terms in the modified qubit Hamiltonian.
@@ IV-C Constant-ancilla method for other mappings
Our method for constructing constant-ancilla fermion-qubit mappings in Section IV-B begins by using the Jordan–Wigner transformation to generate the base ancilla-free Hamiltonian in (39). If we had instead used a different linear encoding, we predict that it would generally be computationally hard to devise an effective set of commuting Pauli strings $P_{0},P_{1},\dots,P_{p-1}$ for incremental ancilla addition because number, hopping and interaction terms map to Pauli strings with less obvious nonlocal structure than the $Z_{i+1}\dots Z_{j-1}$ factors of the Jordan–Wigner transformations. In modifying the Hamiltonian terms from $H_{q}(\sigma)$ to $H_{q}^{\prime}$ with the addition of the $k$th ancilla qubit, the protocol would need to search for Pauli operators $P_{k}=X_{\mathcal{I}_{k}}Y_{\mathcal{J}_{k}}Z_{\mathcal{K}_{k}}$ such that multiplying the terms of $H_{q}(\sigma)$ by $P_{k}$ reduces the total or maximum Pauli weight. As discussed elsewhere [38, 11] a new requirement emerges that $P_{j},P_{k}]=0$ for all $j\in[k]$ in order to satisfy statement 2. Again, the reason this was straightforward for the Jordan–Wigner case was because we knew a priori that the problematic parts of the hopping terms would consist only of $Z$ matrices. Nevertheless, we see this as a worthwhile line of inquiry because of the potential for other mappings to reduce the weight of qubit Hamiltonians from the outset.
TABLE I: A comparison of the Pauli weight for Jordan–Wigner, Bravyi–Kitaev, Parity Basis and Ternary Tree mappings of fermionic systems with Hamiltonian graphs that are square grids of size $\times3$ to $\times15$. Orderings are optimized for total Pauli weight, max Pauli weight, and total Pauli weight assuming real valued coefficients and the absence of operator terms, corresponding to the Fermi–Hubbard model.
@@
Total Pauli weight $\mathcal{D}=\text{Num}+\text{ReHop}$ + $\text{ ImHop}+\text{Inter}$ BK 304 635 1107 1712 2473 3331 4467 5741 7127 8850 10438 12595 14522
@@
$T$ 313 628 1080 1676 2375 3237 4303 5473 6799 8342 9853 11844 13942
Maximum Pauli weight $\mathcal{D}=\max(\text{Num},$ $\text{ReHop},\text{ImHop},\text{Inter})$ BK 5 7 9 9 11 11 12 13 13 13 14 15 15
@@
64-mode system graph JW $T$
@@ V Experiments
[…]
V-A Encoding comparison with optimized fermionic order
In this section, we examine the difference in the total and maximum Pauli weights yielded by the four primary encodings. In Table I, we compare the results of the Jordan–Wigner, Bravyi–Kitaev, Parity Basis and Ternary Tree transformations applied to a series of square grids of increasing size. We show results for three different optimization objectives: total Pauli weight, maximum Pauli weight, and total Pauli weight for a model with only real-valued coefficients and no operator terms, which corresponds to the popular Fermi–Hubbard model [13]. Here, we see that Jordan–Wigner is the preferred encoding for smaller grids, allowing orderings which produce a small maximum and total Pauli weight compared to other methods. However, for all but the smallest input, Ternary Tree remains comparable, and eventually emerges as the superior ordering. Bravy–Kitaev’s better scaling similarly allows it to outperform Jordan–Wigner for larger grids. Within the Jordan–Wigner orderings, we recover the optimal Mitchison–Durbin ordering [28] for square lattices of size $\times6$ and larger, agreeing with the analytical result in [11].
@@ V-B Jordan–Wigner transformations with optimal ancilla use
[…]
@@
64-mode system graph JW $T$ JW + 10 ancillas
@@ VI Conclusion
We have introduced a cost model for the total and maximum Pauli weight of the output of linear encodings, a type of ancilla-free fermion-qubit mapping, and have found optimized fermionic orderings for several popular transformations using quadratic assignment. Empirically, these optimized orderings are able to recover known optimal solutions in the case of the Jordan–Wigner mapping. Moreover, we find an optimal strategy to cancel out disjoint strings of Pauli-$Z$ gates, utilizing ancilla qubits to reduce the Pauli weight of qubit Hamiltonians associated with the Jordan–Wigner transformation. According to our results, the latter strategy requires only of 10 ancilla qubits to reduce Hamiltonian Pauli weight by as much as 67%, outperforming the optimized ancilla-free Ternary Tree transformation. Future work might expand our approach to more complex cost functions that incorporate gate depth and circuit compilation, and generalise our incremental ancilla method to mappings beyond the Jordan–Wigner transformation.